OpenStudy (anonymous):

Solve The Equations Below ; 1) 3x+9=5 2) 8z-1=11

OpenStudy (anonymous):

x=-4/5 z=3/2

jimthompson5910 (jim_thompson5910):

3x+9=5 3x=5-9 3x=-4 x=-4/3

jimthompson5910 (jim_thompson5910):

8z-1=11 8z=11+1 8z=12 z=12/8 z=3/2

OpenStudy (anonymous):

my mistake that sud have been -4/3

OpenStudy (anonymous):

Thankss ' (: